Harry’s Carryout Stores has eight locations. The firm wishes to expand by two more stores and needs a bank loan to do this. Mr. Wilson, the banker, will finance construction if the firm can present an acceptable three-month financial plan for January through March. The following are actual and forecast sales figures:

Actual Forecast Additional Information
November $300,000 January $380,000 April forecast $390,000
December 320,000 February 420,000
March 400,000

Of the firm’s sales, 50 percent are for cash and the remaining 50 percent are on credit. Of credit sales, 35 percent are paid in the month after sale and 65 percent are paid in the second month after the sale. Materials cost 30 percent of sales and are purchased and received each month in an amount sufficient to cover the following month’s expected sales. Materials are paid for in the month after they are received. Labor expense is 40 percent of sales and is paid for in the month of sales. Selling and administrative expense is 15 percent of sales and is paid in the month of sales. Overhead expense is $30,500 in cash per month.

Depreciation expense is $10,500 per month. Taxes of $8,500 will be paid in January, and dividends of $4,500 will be paid in March. Cash at the beginning of January is $90,000, and the minimum desired cash balance is $85,000.

a. Prepare a schedule of monthly cash receipts for January, February, and March.
  

b. Prepare a schedule of monthly cash payments for January, February, and March.
  


c. Prepare a monthly cash budget with borrowings and repayments for January, February, and March. (Negative amounts should be indicated by a minus sign. Assume the January beginning loan balance is $0.)
Answer #1

Part (a)

Please see the table below. Please be guided by the second column titled “Linkage” to understand the mathematics. The last row highlighted in yellow is your answer. Figures in parenthesis, if any, mean negative values. All financials are in $.

Month Linkage November December January February March April
Sales A $300,000 320,000 $380,000 420,000 400,000 $390,000
Cash sales B = 50% x A $150,000 $160,000 $190,000 $210,000 $200,000 $195,000
Credit Sales C = 50% x A $150,000 $160,000 $190,000 $210,000 $200,000 $195,000
Collection from credit sales D = 35% x C(m-1) + 65% x C(m - 2)* $52,500 $153,500 $170,500 $197,000 $206,500
Monthly cash receipts E = B + D $343,500 $380,500 $397,000

* Collection from credit sales = 35% x Credit sales of last month + 65% x Credit sales two months back. To explain, the figure under the head of January = 153,500 = 35% x 160,000 + 65% x 150,000

Part (b)

Cash payments Linkage November December January February March April
Purchases F = A x 30%                96,000             114,000           126,000            120,000          117,000                   -  
Payment on account of purchases G = shift F above by 1 month               96,000           114,000            126,000          120,000        117,000
Labor expenses payment H = 40% x A              120,000             128,000           152,000            168,000          160,000        156,000
Selling & admin expenses payment I = 15% x A                45,000               48,000             57,000              63,000             60,000          58,500
Overhead expenses J                30,500               30,500             30,500              30,500             30,500          30,500
Taxes K                8,500
Dividends L               4,500
Monthly cash payments M = G + H + I + J + K           362,000            387,500          375,000

Part (c)

Month Linkage November December January February March
Cash receipts A $343,500 $380,500 $397,000
Cash payments M           362,000            387,500          375,000
Opening balance in cash N $90,000 $85,000 $85,000
Likely closing balance without any external capital raise O = N + A - M             71,500              78,000          107,000
Minimum Desired closing balance P             85,000              85,000             85,000
Surplus / (Shortfall) Q = O - P            (13,500)               (7,000)             22,000
Fresh borrowing R = max (-Q, 0)             13,500                7,000                     -  
Repayment S = max (Q , 0)             22,000
Actual closing balance T = O + R - S $85,000 $85,000 $85,000